### v1

# The new observations about the parameter-dependent Schrödinger-Poisson system

Abstract.
In this paper, we study the existence results of solutions for the following Schrödinger-Poisson system involving different potentials:
$\begin{cases}-\Delta u+V(x)u-\lambda\phi u=f(u)&\quad\text{in}~\mathbb{R}^{3},\\ -\Delta\phi=u^{2}&\quad\text{in}~\mathbb{R}^{3}.\end{cases}$
We first consider the case that the potential $V$ is positive and radial so that the mountain pass theorem could be implied. The other case is that the potential $V$ is coercive and sign-changing, which means that the Schrödinger operator $-\Delta+V$ is allowed to be indefinite. To deal with this more difficult case, by a local linking argument and Morse theory, the system has a nontrivial solution. Furthermore, we also show the asymptotical behavior result of this solution. Additionally, the proofs rely on new observations regarding the solutions of the Poisson equation. As a main novelty with respect to corresponding results in [21, 12, 20], we only assume that $f$ satisfies the super-linear growth condition at the origin. We believe that the methodology developed here can be adapted to study related problems concerning the existence of solutions for Schrödinger-Poisson system.
Keywords: Schrödinger-Poisson system; Perturbation problems; Mountain pass theorem; Local linking; Non-trivial solutions.
Mathematics Subject Classification 2020: 35A15, 35B38, 35J50
1. Introduction and Main Results
This paper is concerning about the existence of the nontrivial solutions of the nonlinear Schrödinger-Poisson system in the following
(1.1) $\begin{cases}-\Delta u+V(x)u-\lambda\phi u=f(u)&\quad\text{in }\mathbb{R}^{3},\\ -\Delta\phi=u^{2}&\quad\text{in }\mathbb{R}^{3},\end{cases}$
where $V$ and $f$ are given functions, $\lambda$ is a positive constant. To find the mountain pass solution, we only need the following assumptions:
${\bf(V)}$: $V(x)\in C(\mathbb{R}^{3},\mathbb{R})$ is radial, i.e., $V(x)=V(|x|)$, $\inf_{\mathbb{R}^{3}}V(x)\geq V_{0}>0$.
${\bf(f)}$: $f(t)\in C(\mathbb{R},\mathbb{R})$ and $\lim\limits_{t\rightarrow 0}\frac{f(t)}{t}=0.$
System (1.1) arises when we focus on standing waves $\psi(t,x)=e^{-i\omega t/\sqrt{2m}}u(x)$ for the time-dependent Schrödinger equations
$i\sqrt{2m}\frac{\partial\psi}{\partial t}=-\Delta\psi+U(x)\psi-\lambda\phi\psi-g(|\psi|)\psi,\ (t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{3}$
with potential $\phi$ depending on the wave function $\psi$ via the Poisson equation
$\Delta\phi=|\psi|^{2}\quad\text{in}\ \mathbb{R}^{3}.$
That is why we call system (1.1) the stationary Schrödinger Poisson systems. Remark that in system (1.1),
$V(x)=U(x)-\omega\ \text{and}\ f(t)=g(|t|)t.$
The Schrödinger-Poisson system (1.1) has been extensively studied, especially in case of $\lambda<0$. In recent years, there has been an increasing attention to the existence of positive solutions, ground states, sign-changing solutions and multiplicity of solutions related to the system (1.1). Ruiz [23] focused on the following problem
$\begin{cases}-\Delta u+V(x)u-\lambda\phi u=|u|^{p-2}u&\quad\text{in }\mathbb{R}^{3},\\ -\Delta\phi=u^{2}&\quad\text{in }\mathbb{R}^{3},\end{cases}$
and gave the existence and nonexistence results, depending on the parameters $p\in(2,6)$ and $\lambda<0$. In particular, if $\lambda\leq-\frac{1}{4}$, the author showed that $p=3$ is a critical value for the existence of positive solutions. By using the concentration compactness principle, Azzollini and Pomponio in [3] proved the existence of a ground state solution of system (1.1) when $f(u)=|u|^{p-2}u$ and $p\in(3,6)$. Here, we would also like to mention the papers [19, 8, 17, 21, 24, 22, 18, 23, 9, 6, 2, 3, 1].
From the above references, it can be seen that when $\lambda<0$, the existence of solutions to system (1.1) is closely related to the growth of the nonlinear term (i.e., the range of value of $p$). However, when the parameter $\lambda>0$, the situation changed, for $V(x)=V_{0}>0$, Mugnai in [20] proved that there exist infinitely many triples $(\lambda,u,\phi)\in\mathbb{R}^{+}\times H^{1}(\mathbb{R}^{3})\times D^{1,2}(\mathbb{R}^{3})$ solving system (1.1). Here $\lambda$ is a Lagrange multiplier and is part of the solutions obtained, the nonlinearity $f$ satisfies the following conditions:
($\tilde{f}_{1}$): $f\in C(\mathbb{R},\mathbb{R})$ and $f(0)=0$;
($\tilde{f}_{2}$): there are two positive constants $C_{1}$, $C_{2}$ such that
$|f(t)|\leq C_{1}|t|^{p-1}+C_{2}|t|^{q-1}\quad\text{ with}\ \ 2<p<q<6;$
($\tilde{f}_{3}$): there is a positive constant $\mu>2$, such that $0\leq tf(t)\leq\mu F(t)$ with
$F(t)=\int_{0}^{t}f(s)\text{d}s;$
($\tilde{f}_{4}$): $f(t)$ is an odd function.
For $\lambda>0$ sufficiently large, Jeong and Seok in [12] proved the existence of a radial solution under the conditions $V(x)=V_{0}>0$, $({\bf f})$, and the following assumption:
[…]
As a consequence, this generalized the conditions in Mugnai [20].
The method in [12] relies on the non-degeneracy of the solutions of the limiting equation to prove the boundedness of the Palais Smale (PS) sequence of the energy functional corresponding to system (1.1). As a result, it removes the Ambrosetti-Rabinowitz $(AR)$ condition in Mugnai [20], retaining only the growth condition. This is to ensure that the energy functional corresponding to the equation can be defined in a Hilbert space.
A natural question then arises: Can the conditions at infinity for the nonlinear term $f$ be removed, i.e. ${\bf(\tilde{f})}$? This is one of the major innovative points of this paper. Another significant innovation is that our approach is completely different from that in [12]. We have observed an additional property of the solutions of the Poisson equation. By leveraging this property, we are able to prove the boundedness of the $(PS)$ sequence of the energy functional corresponding to system (1.1) without relying on the non-degeneracy of the original problem.
Before stating our the first main result, we will introduce some preliminaries. Consider the Hilbert space as a working space,
$H^{1}_{r}(\mathbb{R}^{3})=\{u\in H^{1}(\mathbb{R}^{3}):u(x)=u(|x|)\}$
with the norm
$\|u\|_{H^{1}_{r}(\mathbb{R}^{3})}:=\left(\int_{\mathbb{R}^{3}}\left(|\nabla u|^{2}+V(x)u^{2}\right)\text{d}x\right)^{\frac{1}{2}}.$
It is well-known that $H^{1}_{r}(\mathbb{R}^{3})$ is compactly embedded in $L^{p}(\mathbb{R}^{3})(2<p<6)$.
Let $\lambda=\frac{1}{\epsilon^{2}}$ with $\epsilon>0$, by making a transformation
$\bar{u}=\frac{u}{\epsilon}$
with abuse of the notations (still labeled $\bar{u}$ by $u$) the system (1.1) is equivalent to the following
(1.2) $\begin{cases}-\Delta u+V(x)u-\phi u=\frac{1}{\epsilon}f(\epsilon u)&\quad\text{in }\mathbb{R}^{3},\\ -\Delta\phi=u^{2}&\quad\text{in }\mathbb{R}^{3}.\end{cases}$
[…]
(1.3) $\phi_{u}(x)=\frac{1}{4\pi}\int_{\mathbb{R}^{3}}\frac{u^{2}(y)}{|x-y|}\text{d}y.$
By [12, Proposition 7], one has
(1.4) $\int_{\mathbb{R}^{3}}|\nabla\phi_{u}|^{2}\text{d}x=\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\text{d}x\leq C\|u\|_{H^{1}_{r}(\mathbb{R}^{3})}^{4}.$
Here the subscript $u$ of $\phi$ denotes the solution of system (1.2) corresponding to $u^{2}$. The constant $C$ here and subsequently in the rest of this paper may change from line to line as long as what these constants depend on is clear.
Now, we present our first result.
Theorem 1.
Under assumptions ${\bf(V)}$ and ${\bf(f)}$. Thne, there is a constant $\lambda^{*}>0$ such that if $\lambda\geq\lambda^{*}$, system (1.1) admits a non-trivial weak solution $u$ in the following sense
$\int_{\mathbb{R}^{3}}\left(\nabla u\nabla\psi+V(x)u\psi\right)\text{d}x-\lambda\int_{\mathbb{R}^{3}}\phi u\psi\text{d}x=\int_{\mathbb{R}^{3}}f(u)\psi\text{d}x\ \ ~\text{for~all}~\psi\in C_{0}^{\infty}(\mathbb{R}^{3})$
and
$\int_{\mathbb{R}^{3}}\nabla\phi\nabla\eta\text{d}x=\int_{\mathbb{R}^{3}}u^{2}\eta\text{d}x\ \ ~\text{for~all}~\eta\in C_{0}^{\infty}(\mathbb{R}^{3}).$
Remark 1.
In our approach, the following inequality
$\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\text{d}x=\int_{\mathbb{R}^{3}}|\nabla\phi_{u}|^{2}\text{d}x\geq\int_{\mathbb{R}^{3}}|u|^{3}\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}|\nabla u|^{2}\text{d}x$
is crucial. It is worth emphasizing that this inequality depends solely on the properties of the Poisson equation itself. To the best of our knowledge, few scholars have utilized this inequality to prove the boundedness of the $(PS)$ sequence. To some extent, our work tremendously improves the related results in [20, 12] mentioned above.
We emphasize that in Theorem 1, we only considered the case where the Schrödinger operator $-\Delta+V$ is positive definite. In this situation the zero function $u={\bf 0}$ is a local minimizer of $I$ and one may apply the mountain pass theorem.
However, if we are interested in standing waves $\psi(t,x)=e^{-i\omega t/\sqrt{2m}}u(x)$ with large $\omega$, then $V(x)$ may be negative somewhere and $u={\bf 0}$ is not a local minimizer of $I$. When $\lambda<0$, the corresponding indefinite situation is studied in Liu and Wu [15].
It is interesting to note that unlike many works in this field, the second result in this paper is about the case $\lambda>0$ and the potential $V$ is sign-changing. We present our assumptions on the potential $V$ and the nonlinearity $f$.
${\bf(V_{1})}$: $V\in C(\mathbb{R}^{3})$ is bounded from below and, $\mu(V^{-1}(-\infty,M]))<\infty$ for all $M>0$, where $\mu$ is the Lebesgue measure on $\mathbb{R}^{3}$;
${\bf(V_{2})}$: $0$ is not an eigenvalue of $-\Delta+V(x)$.
${\bf(f)}$: $f\in C(\mathbb{R},\mathbb{R})$ and
$\lim\limits_{t\to 0}\frac{f(t)}{t}=0;$
${\bf(f^{\prime})}$: there exists $\delta_{0}>0$ and for all $|t|\leq\delta_{0}$ such that
$f(t)t\geq 0.$
Inspired by [15, 16], by combining a local linking argument of Li and Willem [13] and infinite-dimensional Morse theorem [7], we will obtain our main result as follows:
Theorem 2.
Suppose that assumptions ${\bf(V_{1})}$, ${\bf(V_{2})}$, ${\bf(f)}$ and ${\bf(f^{\prime})}$ hold, then there is a constant $\lambda^{**}>0$ such that for any $\lambda\in[\lambda^{**},+\infty)$, system (1.1) has a nontrivial solution $u$ in the following sense
$\int_{\mathbb{R}^{3}}\left(\nabla u\nabla\psi+V(x)u\psi\right)\text{d}x-\lambda\int_{\mathbb{R}^{3}}\phi u\psi\text{d}x=\int_{\mathbb{R}^{3}}f(u)\psi\text{d}x\ \ ~\text{for~all}~\psi\in C_{0}^{\infty}(\mathbb{R}^{3})$
and
$\int_{\mathbb{R}^{3}}\nabla\phi\nabla\eta\text{d}x=\int_{\mathbb{R}^{3}}u^{2}\eta\text{d}x\ \ ~\text{for~all}~\eta\in C_{0}^{\infty}(\mathbb{R}^{3}).$
Remark 2.
What we need to note is that Theorem 2 could be understood in another perspective. By assumption ${\bf(V_{1})}$, one choose $m>0$ such that
(1.5) $\tilde{V}(x)=V(x)+m>1$
for all $x\in\mathbb{R}^{3}$. Then the system (1.1) is equivalent to the following
$\begin{cases}-\Delta u+\tilde{V}(x)u-\lambda\phi u=\tilde{f}(u)&\quad\text{in }\mathbb{R}^{3},\\ -\Delta\phi=u^{2}&\quad\text{in }\mathbb{R}^{3},\end{cases}$
with $\tilde{f}(u)=f(u)+mu$. The straightforward computation gives the potential $\tilde{V}(x)$ is positive and coercive, and $\tilde{f}(u)$ satisfies local linear growth:
$\tilde{f}(t)\in C(\mathbb{R},\mathbb{R})\ \text{and}\ \lim\limits_{t\rightarrow 0}\frac{\tilde{f}(t)}{t}=m>0.$
Thus the essential difference between Theorem 1 and Theorem 2 lies in the different states of the nonlinearity at the origin.
Remark 3.
The innovations of this paper are as follows:
$(1)$ In terms of results, through Theorem 1 and Theorem 2, we show that when the parameter is sufficiently large, the existence of solutions to the system depends only on the state of the nonlinear term at the origin.
$(2)$ In terms of methods, we propose a completely new modification method. Additionally, by means of new observations on the solutions to Poisson’s equation, we obtain the boundedness of the $(PS)$ sequence of the modified functional. To the best of our knowledge, this phenomenon is also a new discovery.
We also study the asymptotical behavior of non-trivial solution $u^{\lambda}$ with respect to $\lambda$.
Theorem 3.
Under assumptions ${\bf(V)}$ and ${\bf(f)}$. Let $u^{\lambda}$ be a non-trivial weak solutions of system (1.1), then as $\lambda\to+\infty$,
$\begin{split}u^{\lambda}&\to 0~\text{strongly~in}~H^{1}_{r}(\mathbb{R}^{3}),\\ \sqrt{\lambda}u^{\lambda}&\to u_{0}~\text{strongly~in}~H^{1}_{r}(\mathbb{R}^{3}),\end{split}$
where $u_{0}$ is a non-trivial weak solution of the following Choquard equation
$\Delta u+V(x)u=\frac{1}{4\pi}\int_{\mathbb{R}^{3}}\frac{u^{2}(y)}{|x-y|}\text{d}yu\ \ ~\text{in}~\mathbb{R}^{3}.$
The paper is organized as follows. In Section 2, we propose a new modified problem. From this new modification method, it can be seen that there is a close relationship between the solution of this modified problem and the solution of the original problem. In Section 3, we prove that the modified functional has critical points, and when the parameter is sufficiently large, the critical point is exactly the critical point of the original problem. In Section 4, we prove that similar results in the case of sign-changing potentials. In Section 5, we give the proof of the asymptotical behavior of non-trival solutions.
2. The modified problem
Since the condition ${\bf(f)}$ only involves the information at the origin, a truncation of $F$ is introduced such that the corresponding functional is well-defined in the working space. It is worth emphasizing that this modification method is new and specially constructed based on the system (1.1).
Write
$d(t)\begin{cases}=1,&\quad\text{if }|t|<1,\\ \in(0,1),&\quad\text{if }1\leq|t|\leq 2,\\ =0,&\quad\text{if }|t|>2,\end{cases}$
and $d(t)\in C_{0}^{\infty}(\mathbb{R})$ satisfying $-C(d)\leq d^{\prime}(t)t\leq 0$, where $C(d)$ is a positive constant. The truncation about $F$ is defined as
(2.1) $\tilde{F}(t)=d\left(\frac{t}{\sqrt{\epsilon}}\right)F(t)+\left(1-d\left(\frac{t}{\sqrt{\epsilon}}\right)\right)|t|^{3},$
where $\epsilon\in(0,1)$ is a constant. Then the directly calculation gives that
$\begin{split}|\tilde{F}^{\prime}(t)|&\leq\left|d^{\prime}(\frac{t}{\sqrt{\epsilon}})F(t)\frac{1}{\sqrt{\epsilon}}\right|+\left|d(\frac{t}{\sqrt{\epsilon}})f(t)\right|+\left|d^{\prime}(\frac{t}{\sqrt{\epsilon}})\frac{1}{\sqrt{\epsilon}}|t|^{3}\right|+\left|3(1-d(\frac{t}{\sqrt{\epsilon}}))t^{2}\right|\\ &=\left|d^{\prime}(\frac{t}{\sqrt{\epsilon}})\frac{F(t)}{t}\frac{t}{\sqrt{\epsilon}}\right|+\left|d(\frac{t}{\sqrt{\epsilon}})f(t)\right|+\left|d^{\prime}(\frac{t}{\sqrt{\epsilon}})\frac{t}{\sqrt{\epsilon}}t^{2}\right|+\left|3(1-d(\frac{t}{\sqrt{\epsilon}}))t^{2}\right|\\ &\leq\left|d^{\prime}(\frac{t}{\sqrt{\epsilon}})\frac{t}{\sqrt{\epsilon}}\right|\left|\frac{F(t)}{t}\right|+\left|d(\frac{t}{\sqrt{\epsilon}})f(t)\right|+\left|d^{\prime}(\frac{t}{\sqrt{\epsilon}})\frac{t}{\sqrt{\epsilon}}\right|t^{2}+\left|3(1-d(\frac{t}{\sqrt{\epsilon}}))t^{2}\right|.\end{split}$
For any $\delta>0$, it follows from $-C(d)\leq d^{\prime}(t)t\leq 0$ and ${\bf(f)}$ that there is a constant $C_{\delta}$ depending on $\delta$ such that
(2.2) $|\tilde{F}(t)|\leq\delta t^{2}+C_{\delta}|t|^{3}\quad\text{and}\quad|\tilde{F}^{\prime}(t)|\leq\delta|t|+C_{\delta}|t|^{2},$
where $C_{\delta}$ is a constant independent of $\epsilon$.
According to system (2.1), the following modified problem is considered instead of system (1.1)
(2.3) $\begin{cases}-\Delta u+V(x)u-\phi u=\frac{1}{\epsilon}\tilde{F}^{\prime}(\epsilon u)&\quad\text{in }\mathbb{R}^{3},\\ -\Delta\phi=u^{2}&\quad\text{in }\mathbb{R}^{3}.\end{cases}$
Therefore, the corresponding functional is defined as
(2.4) $\text{$\displaystyle I_{\epsilon}(u)$}=\frac{1}{2}\int_{\mathbb{R}^{3}}\left(|\nabla u|^{2}+V(x)u^{2}\right)\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\text{d}x-\frac{1}{\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}(\epsilon u)\text{d}x.$
The directly calculation gives that $I_{\epsilon}(u)$ is a $C^{1}$ functional on $H^{1}_{r}(\mathbb{R}^{3})$ and for $u,v\in H^{1}_{r}(\mathbb{R}^{3})$, one has
$\langle I_{\epsilon}^{\prime}(u),v\rangle=\int_{\mathbb{R}^{3}}\left(\nabla u\nabla v+V(x)uv\right)\text{d}x-\int_{\mathbb{R}^{3}}\phi_{u}uv\text{d}x-\frac{1}{\epsilon}\int_{\mathbb{R}^{3}}\tilde{F}^{\prime}(\epsilon u)v\text{d}x.$
3. The existence of parameter-dependent Schrödinger-Poisson system
In this section, the non-trivial solution of systme (1.1) is obtained and give the proof of Theorem 1. Firstly, the mountain pass theorem is applied to prove the existence of critical points for the modified functional $I_{\epsilon}$. Secondly, the truncation of $\tilde{F}$ is removed when $\epsilon>0$ is sufficiently small by Moser iteration. Finally, the mountain pass critical point of the modified functional $I_{\epsilon}$ is actually the solution to the original system (1.1).
Lemma 1.
$I_{\epsilon}(u)$ possesses the following properties:
($i$) There are constants $\rho>0$ and $\beta>0$ independent of $\epsilon$ such that
$I_{\epsilon}(u)\geq\beta>0\text{ for $\displaystyle u\in S_{\rho}=\{u:\|u\|_{H^{1}_{r}(\mathbb{R}^{3})}=\rho\}$}.$
($ii$) There exists a function $\zeta\in H^{1}_{r}(\mathbb{R}^{3})$ such that $\|\zeta\|_{H^{1}_{r}(\mathbb{R}^{3})}>\rho$ and $I_{\epsilon}(\zeta)<0$.
Proof.
It follows from (1.4) and (2.2) that
[…]
Choosing $\delta>0$ small, one gives
$\begin{split}I_{\epsilon}(u)\geq\frac{1}{3}\|u\|^{2}_{H^{1}_{r}(\mathbb{R}^{3})}-\frac{C}{4}\|u\|^{4}_{H^{1}_{r}(\mathbb{R}^{3})}-C\|u\|^{3}_{H^{1}_{r}(\mathbb{R}^{3})}\end{split},$
where the Sobolev inequality has been used. This implies ($i$).
For ($ii$), take a $h\in C_{0}^{\infty}(\mathbb{R}^{3})$, $h\neq 0$ and let $s$ be a constant large enough. Substituting $sh$ in $I_{\epsilon}$ yields
$\begin{split}I_{\epsilon}(sh)&\leq\frac{s^{2}}{2}\|h\|^{2}_{H^{1}_{r}(\mathbb{R}^{3})}-\frac{s^{4}}{4}\int_{\mathbb{R}^{3}}\phi_{h}h^{2}\text{d}x+\frac{1}{\epsilon^{2}}\int_{\mathbb{R}^{3}}\left(\delta\epsilon^{2}s^{2}h^{2}+C_{\delta}\epsilon^{3}s^{3}|h|^{3}\right)\text{d}x\\ &\leq\frac{s^{2}}{2}\|h\|^{2}_{H^{1}_{r}(\mathbb{R}^{3})}-\frac{s^{4}}{4}\int_{\mathbb{R}^{3}}\phi_{h}h^{2}\text{d}x+\int_{\mathbb{R}^{3}}\left(\delta s^{2}h^{2}+C_{\delta}s^{3}|h|^{3}\right)\text{d}x.\end{split}$
Thanks to $\int_{\mathbb{R}^{3}}\phi_{h}h^{2}\text{d}x\geq 0$, then $I_{\epsilon}(sh)\rightarrow$-$\infty$ as $s\rightarrow+\infty.$ Hence, the lemma is completed with $\zeta=sh$. ∎
According to Lemma 1, it follows from the mountain pass theorem in Willem [25] that there is a sequence $\{u_{n}\}_{n=1}^{\infty}\subset H^{1}_{r}(\mathbb{R}^{3})$ such that
(3.1) $I_{\epsilon}(u_{n})\rightarrow c\quad\text{and}\quad I_{\epsilon}^{\prime}(u_{n})\rightarrow 0\ \text{ as }n\rightarrow+\infty$
with $c$ be a positive constant.
Lemma 2.
Let $\{u_{n}\}_{n=1}^{\infty}\subset H^{1}_{r}(\mathbb{R}^{3})$ satisfies (3.1). There exists a constant $\epsilon_{*}\in(0,1)$ such that if $0<\epsilon<\epsilon_{*}$, then $u_{n}\rightarrow u\text{ strongly in }H^{1}_{r}(\mathbb{R}^{3})$ with $u\in H^{1}_{r}(\mathbb{R}^{3})$.
Proof.
Let $\mu$ be a positive constant to be determined later. It follows from (2.2) that
(3.2) $\begin{split}&\quad I_{\epsilon}(u_{n})-\frac{1}{\mu}\langle I_{\epsilon}^{\prime}(u_{n}),u_{n}\rangle\\ &=\left(\frac{1}{2}-\frac{1}{\mu}\right)\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+V(x)u_{n}^{2}\right)\text{d}x+\left(\frac{1}{\mu}-\frac{1}{4}\right)\int_{\mathbb{R}^{3}}\phi_{u_{n}}u_{n}^{2}\text{d}x-\frac{1}{\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}(\epsilon u_{n})\text{d}x\\ &\quad\quad-\frac{1}{\mu\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}^{\prime}(\epsilon u_{n})\epsilon u_{n}\text{d}x\\ &\geq\left(\frac{1}{2}-\frac{1}{\mu}\right)\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+V(x)u_{n}^{2}\right)\text{d}x+\left(\frac{1}{\mu}-\frac{1}{4}\right)\int_{\mathbb{R}^{3}}\phi_{u_{n}}u_{n}^{2}\text{d}x\\ &\quad\quad-\left(1+\frac{1}{\mu}\right)\int_{\mathbb{R}^{3}}\left(\delta|u_{n}|^{2}+\epsilon C_{\delta}|u_{n}|^{3}\right)\text{d}x.\end{split}$
[…]
(3.3) $\int_{\mathbb{R}^{3}}|u_{n}|^{3}\text{d}x=\int_{\mathbb{R}^{3}}{\nabla\phi_{u_{n}}}\nabla(|u_{n}|)\text{d}x\leq\int_{\mathbb{R}^{3}}|\nabla\phi_{u_{n}}|^{2}\text{d}x+\frac{1}{4}\int_{\mathbb{R}^{3}}|\nabla u_{n}|^{2}\text{d}x.$
Combining (1.4) and (3.3) yields
(3.4) $\int_{\mathbb{R}^{3}}\phi_{u_{n}}u_{n}^{2}\text{d}x=\int_{\mathbb{R}^{3}}|\nabla\phi_{u_{n}}|^{2}\text{d}x\geq\int_{\mathbb{R}^{3}}|u_{n}|^{3}\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}|\nabla u_{n}|^{2}\text{d}x.$
It follows from (3.2) and (3.4) that
$\begin{split}&\quad I_{\epsilon}(u_{n})-\frac{1}{\mu}\langle I_{\epsilon}^{\prime}(u_{n}),u_{n}\rangle\\ &\geq\left(\left(\frac{1}{2}-\frac{1}{\mu}\right)-\frac{1}{4}\left(\frac{1}{\mu}-\frac{1}{4}\right)-\delta\left(1+\frac{1}{\mu}\right)\right)\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+V(x)u_{n}^{2}\right)\text{d}x\\ &\quad\quad+\left(\frac{1}{\mu}-\frac{1}{4}-C_{\delta}\epsilon\right)\int_{\mathbb{R}^{3}}|u_{n}|^{3}\text{d}x.\end{split}$
Let $\mu=3$ and $\delta=\frac{1}{64}$. Then, there exists a constant $\epsilon_{*}>0$ small enough such that when $0<\epsilon\leq\epsilon_{*}$, one has
$\left(\frac{1}{3}-\frac{1}{4}-C_{\frac{1}{64}}\epsilon\right)\geq 0.$
Consequently,
$I_{\epsilon}(u_{n})-\frac{1}{3}\langle I_{\epsilon}^{\prime}(u_{n}),u_{n}\rangle\geq\frac{1}{8}\|u_{n}\|^{2}_{H^{1}_{r}(\mathbb{R}^{3})}.$
On the other hand, since $\{u_{n}\}_{n=1}^{\infty}$ is a (PS) sequence for $I_{\epsilon}$, it implies
$c+o_{n}(1)+o_{n}(1)\|u_{n}\|_{H^{1}_{r}(\mathbb{R}^{3})}\geq I_{\epsilon}(u_{n})-\frac{1}{3}\langle I_{\epsilon}^{\prime}(u_{n}),u_{n}\rangle.$
Combining the above two estimates yields
$\|u_{n}\|_{H^{1}_{r}(\mathbb{R}^{3})}<+\infty.$
Then there is a subsequence of $\{u_{n}\}_{n=1}^{\infty}$ (still labeled by $u_{n}$) and $u\in H^{1}_{r}(\mathbb{R}^{3})$ such that
[…]
The straightforward computations give
(3.6) $\begin{split}&\quad\langle I_{\epsilon}^{\prime}(u_{n})-I_{\epsilon}^{\prime}(u),u_{n}-u\rangle\\ &=\|u_{n}-u\|^{2}_{H^{1}_{r}(\mathbb{R}^{3})}-\int_{\mathbb{R}^{3}}(\phi_{u_{n}}u_{n}-\phi_{u}u)(u_{n}-u)\text{d}x\\ &\quad\quad-\frac{1}{\epsilon}\int_{\mathbb{R}^{3}}(\tilde{F}^{\prime}(\epsilon u_{n})-\tilde{F}^{\prime}(\epsilon u))(u_{n}-u)\text{d}x.\end{split}$
In terms of (3.1), one has
(3.7) $\frac{1}{\epsilon}\int_{\mathbb{R}^{3}}\tilde{F}^{\prime}(\epsilon u)(u_{n}-u)\text{d}x\rightarrow 0\ \ \text{ as }n\rightarrow+\infty.$
Furthermore, due to (2.2), one has
$\begin{split}&\quad\frac{1}{\epsilon}\int_{\mathbb{R}^{3}}\left|\tilde{F}^{\prime}(\epsilon u_{n})(u_{n}-u)\right|\text{d}x\\ &\leq\int_{\mathbb{R}^{3}}\left(\delta|u_{n}||u_{n}-u|+\epsilon C_{\delta}u_{n}^{2}|u_{n}-u|\right)\text{d}x\\ &\leq\delta\|u_{n}\|_{L^{2}(\mathbb{R}^{3})}\|u_{n}-u\|_{L^{2}(\mathbb{R}^{3})}+\epsilon C_{\delta}\left(\int_{\mathbb{R}^{3}}|u_{n}|^{3}\text{d}x\right)^{\frac{2}{3}}\left(\int_{\mathbb{R}^{3}}|u_{n}-u|^{3}\text{d}x\right)^{\frac{1}{3}}.\end{split}$
This combining with (3.7) gives that
(3.8) $\left|\frac{1}{\epsilon}\int_{\mathbb{R}^{3}}(\tilde{F}^{\prime}(\epsilon u_{n})-\tilde{F}^{\prime}(\epsilon u))(u_{n}-u)\text{d}x\right|\leq C\delta\ \ \text{as}~n\rightarrow+\infty.$
It follows from [12, Lemma 1] that
$\begin{split}&\quad\left|\int_{\mathbb{R}^{3}}(\phi_{u_{n}}u_{n}-\phi_{u}u)(u_{n}-u)\text{d}x\right|\\ &\leq\int_{\mathbb{R}^{3}}\left(\left|\phi_{u_{n}}u_{n}(u_{n}-u)\right|+\left|\phi_{u}u(u_{n}-u)\right|\right)\text{d}x\\ &\leq\|\phi_{u_{n}}\|_{L^{6}{(\mathbb{R}^{3})}}\|{u_{n}}\|_{L^{12/5}{(\mathbb{R}^{3})}}\|{u_{n}-u}\|_{L^{12/5}{(\mathbb{R}^{3})}}+\|\phi_{u}\|_{L^{6}{(\mathbb{R}^{3})}}\|{u}\|_{L^{12/5}{(\mathbb{R}^{3})}}\|{u_{n}-u}\|_{L^{12/5}{(\mathbb{R}^{3})}}.\end{split}$
Note that $\phi_{u_{n}}\rightarrow\phi_{u}$ in $L^{6}(\mathbb{R}^{3})$ and $u_{n}\rightarrow u$ in $L^{12/5}{(\mathbb{R}^{3})}$. Then
(3.9) $\left|\int_{\mathbb{R}^{3}}(\phi_{u_{n}}u_{n}-\phi_{u}u)(u_{n}-u)\text{d}x\right|\rightarrow 0\ \ \text{ as }n\rightarrow 0.$
Hence, substituting (3.8) and (3.9) into (3.6) yields
$\|u_{n}-u\|^{2}_{H^{1}_{r}(\mathbb{R}^{3})}\leq o_{n}(1)+C\delta.$
Since $\delta$ is arbitrary, one has
$\|u_{n}-u\|_{H^{1}_{r}(\mathbb{R}^{3})}=0\ \ \text{as }n\rightarrow+\infty.$
This completes the lemma. ∎
By the Lemma 2, there is a $u_{\epsilon}\in H^{1}_{r}(\mathbb{R}^{3})$ such that $I_{\epsilon}(u_{\epsilon})=c_{\epsilon}$ and $I_{\epsilon}^{\prime}(u_{\epsilon})=0$. Therefore, $u_{\epsilon}$ is the weak solution of system (2.3). The following lemma enables us to remove the truncation (2.1).
Lemma 3.
For any $R>1$, there exists $\epsilon^{\prime}(R)>0$ such that for $0<\epsilon<\epsilon^{\prime}(R)$, let $u_{\epsilon}$ be the weak solution of system (2.3) obtained in Lemma 2. Then there exists a positive constant $m_{0}$ independent of $\epsilon$ and $R$ such that
(3.10) $\|u_{\epsilon}\|_{L^{\infty}(\mathbb{R}^{3})}\leq m_{0}.$
Proof.
For any $0<\epsilon<\epsilon_{*}(R)$, let $u_{\epsilon}$ satisfy $I_{\epsilon}(u_{\epsilon})=c_{\epsilon}$ and $I_{\epsilon}^{\prime}(u_{\epsilon})=0$. Using the similar method in Lemma 2, it implies
$I_{\epsilon}(u_{\epsilon})=I_{\epsilon}(u_{\epsilon})-\frac{1}{3}\langle I_{\epsilon}^{\prime}(u_{\epsilon}),u_{\epsilon}\rangle\geq\frac{1}{8}\|u_{\epsilon}\|^{2}_{H^{1}_{r}(\mathbb{R}^{3})}.$
It is easy to say that $\{u_{\epsilon}\}$ is a bounded sequence. Therefore, there exists $M_{0}>0$ independent of $\epsilon$ such that for all $\epsilon\in(0,\epsilon_{*}(R))$, it holds
$\left\|u_{\epsilon}\right\|_{H_{r}^{1}\left(\mathbb{R}^{3}\right)}\leq M_{0}.$
For any $\alpha>1$, define
$\varphi=u_{\epsilon}\left|u_{\epsilon,k}\right|^{2\left(\alpha-1\right)}~\text{and}~u_{\epsilon,k}=\begin{cases}u_{\epsilon},~\text{if}~\left|u_{\epsilon}\right|\leq k,\\ k,\ ~~\text{if}~\left|u_{\epsilon}\right|>k.\end{cases}$
Since $u_{\epsilon}$ is a critical point of $I_{\epsilon}\left(u\right)$, one has $\left<I_{\epsilon}^{\prime}\left(u_{\epsilon}\right),\varphi\right>=0$, i.e.,
(3.11) $\int_{\mathbb{R}^{3}}\left(\nabla u_{\epsilon}\nabla\varphi+V(x)u_{\epsilon}\varphi\right)\text{d}x=\int_{\mathbb{R}^{3}}\phi_{u_{\epsilon}}u_{\epsilon}\varphi\text{d}x-\frac{1}{\epsilon}\int_{\mathbb{R}^{3}}\widetilde{f}_{R}\left(\epsilon u_{\epsilon}\right)\varphi\text{d}x.$
For simplicity, we refer to $u_{\epsilon}$ as $u$, $u_{\epsilon,k}$ as $u_{k}$. Substituting $\varphi=u\left|u_{k}\right|^{2\left(\alpha-1\right)}$ into the left side of (3.11) yields
$\quad\int_{\mathbb{R}^{3}}\left|u_{k}\right|^{2\left(\alpha-1\right)}\left|\nabla u\right|^{2}\text{d}x+\int_{\mathbb{R}^{3}}V(x)u^{2}\left|u_{k}\right|^{2\left(\alpha-1\right)}\text{d}x+2\left(\alpha-1\right)\int_{\mathbb{R}^{3}}uu_{k}\left|u_{k}\right|^{2\alpha-4}\nabla u\nabla u_{k}\text{d}x$
$\geq\frac{1}{\alpha^{2}}\int_{\mathbb{R}^{3}}\left|\nabla w_{k}\right|^{2}\text{d}x+\int_{\mathbb{R}^{3}}V(x)w_{k}^{2}\text{d}x$
$\geq\frac{C}{\alpha^{2}}\left(\int_{\mathbb{R}^{3}}w_{k}^{6}\text{d}x\right)^{\frac{1}{3}}+V_{0}\int_{\mathbb{R}^{3}}w_{k}^{2}\text{d}x,$
where $w_{k}=u\left|u_{k}\right|^{\alpha-1}$.
Substituting $\varphi=u\left|u_{k}\right|^{2\left(\alpha-1\right)}$ into the right side of (3.11) yields
$\quad\left|\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\left|u_{k}\right|^{2\left(\alpha-1\right)}\text{d}x-\frac{1}{\epsilon}\int_{\mathbb{R}^{3}}\widetilde{f}_{R}\left(\epsilon u\right)u\left|u_{k}\right|^{2\left(\alpha-1\right)}\text{d}x\right|$
$\leq\left(\int_{\mathbb{R}^{3}}\phi_{u}^{6}\text{d}x\right)^{\frac{1}{6}}\left(\int_{\mathbb{R}^{3}}\left|u^{2}\left|u_{k}\right|^{2\left(\alpha-1\right)}\right|^{\frac{6}{5}}\text{d}x\right)^{\frac{5}{6}}+\frac{1}{\epsilon}\int_{\mathbb{R}^{3}}\left|\widetilde{f}_{R}\left(\epsilon u\right)u\right|\left|u_{k}\right|^{2\left(\alpha-1\right)}\text{d}x$
[…]
$\leq CM^{2}_{0}\left(\int_{\mathbb{R}^{3}}\left|u^{2}\left|u_{k}\right|^{2\left(\alpha-1\right)}\right|^{\frac{6}{5}}\text{d}x\right)^{\frac{5}{6}}+\delta\int_{\mathbb{R}^{3}}u^{2}\left|u_{k}\right|^{2\left(\alpha-1\right)}\text{d}x+C_{\delta,R}\int_{\mathbb{R}^{3}}\epsilon\left|u\right|^{3}\left|u_{k}\right|^{2\left(\alpha-1\right)}\text{d}x.$
Then take $\delta=\frac{V_{0}}{2}$ and $\epsilon(R)=C_{\frac{V_{0}}{2},R}^{-1}$ for $0<\epsilon\leq\epsilon(R)$, one has
$\quad\left|\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\left|u_{k}\right|^{2\left(\alpha-1\right)}\text{d}x-\frac{1}{\epsilon}\int_{\mathbb{R}^{3}}\widetilde{f}_{R}\left(\epsilon u\right)u\left|u_{k}\right|^{2\left(\alpha-1\right)}\text{d}x\right|$
$\leq C_{1}\left(\int_{\mathbb{R}^{3}}\left|w_{k}\right|^{\frac{12}{5}}\text{d}x\right)^{\frac{5}{6}}+\frac{V_{0}}{2}\int_{\mathbb{R}^{3}}w_{k}^{2}\text{d}x+\epsilon C_{\frac{V_{0}}{2},R}\int_{\mathbb{R}^{3}}w_{k}^{2}\left|u\right|\text{d}x$
$\leq C_{1}\left(\int_{\mathbb{R}^{3}}\left|w_{k}\right|^{\frac{12}{5}}\text{d}x\right)^{\frac{5}{6}}+\epsilon C_{\frac{V_{0}}{2},R}\left(\int_{\mathbb{R}^{3}}u^{6}\text{d}x\right)^{\frac{1}{6}}\left(\int_{\mathbb{R}^{3}}|w_{k}|^{\frac{12}{5}}\text{d}x\right)^{\frac{5}{6}}+\frac{V_{0}}{2}\int_{\mathbb{R}^{3}}w_{k}^{2}\text{d}x$
$\leq C_{1}\left(\int_{\mathbb{R}^{3}}\left|w_{k}\right|^{\frac{12}{5}}\text{d}x\right)^{\frac{5}{6}}+\left(\int_{\mathbb{R}^{3}}\left|w_{k}\right|^{\frac{12}{5}}\text{d}x\right)^{\frac{5}{6}}+\frac{V_{0}}{2}\int_{\mathbb{R}^{3}}w_{k}^{2}\text{d}x.$
Combining the two inequalities above yields
$\frac{C}{\alpha^{2}}\left(\int_{\mathbb{R}^{3}}w_{k}^{6}\text{d}x\right)^{\frac{1}{3}}\leq(C_{1}+1)\left(\int_{\mathbb{R}^{3}}\left|w_{k}\right|^{\frac{12}{5}}\text{d}x\right)^{\frac{5}{6}}.$
Then one gets
$\frac{1}{\alpha^{2}}\left\|w_{k}\right\|_{L^{6}(\mathbb{R}^{3})}^{2}\leq\left(C_{1}+C_{2}\right)\left\|w_{k}\right\|_{L^{\frac{12}{5}}(\mathbb{R}^{3})}^{2}.$
Taking $\mu^{*}=\frac{12}{5}\in\left(2,6\right)$ gives
(3.12) $\left\|w_{k}\right\|_{L^{6}(\mathbb{R}^{3})}\leq\alpha C\left\|w_{k}\right\|_{L^{\mu^{\ast}}(\mathbb{R}^{3})}.$
If $u^{\alpha}\in L^{\frac{12}{5}}\left(\mathbb{R}^{3}\right)$ and let $k\to+\infty$ in both side of (3.12), one has
(3.13) $\left\|u\right\|_{L^{6\alpha}(\mathbb{R}^{3})}\leq\alpha^{\frac{1}{\alpha}}C^{\frac{1}{\alpha}}\left\|u\right\|_{L^{\mu^{\ast}\alpha}(\mathbb{R}^{3})}.$
Set $\alpha_{n}=\left(\frac{5}{2}\right)^{n}$, $n=1,2,\cdots$. Clearly, $\alpha_{1}=\frac{5}{2}>1$, $2^{*}\alpha_{n}=\mu^{*}\alpha_{n+1}$, i.e., $6\alpha_{n}=\frac{12}{5}\alpha_{n+1}$ and $\alpha_{n}\to\infty$. Iterating (3.13) with the sequence $\left\{\alpha_{n}\right\}$ gives
(3.14) $\left\|u\right\|_{L^{6\alpha_{n}}(\mathbb{R}^{3})}\leq\alpha_{1}^{\sum_{i=1}^{n}\frac{i}{\alpha_{i}}}C^{\sum_{i=1}^{n}\frac{1}{\alpha_{i}}}\left\|u\right\|_{L^{6}(\mathbb{R}^{3})}~\text{for~all}~n\in\mathbb{N}.$
Note that $\sum_{i=1}^{\infty}\frac{i}{\alpha_{i}}=\frac{10}{9}$ and $\sum_{i=1}^{\infty}\frac{1}{\alpha_{i}}=\frac{2}{3}$. Taking $n\to+\infty$ on both sides of (3.14) gives
$\left\|u\right\|_{L^{\infty}(\mathbb{R}^{3})}\leq\left(\frac{5}{2}\right)^{\frac{10}{9}}C^{\frac{2}{3}}\left\|u\right\|_{L^{6}(\mathbb{R}^{3})}\leq C^{\prime}\left\|u\right\|_{H_{r}^{1}(\mathbb{R}^{3})}.$
Therefore, we get
$\left\|u_{\epsilon}\right\|_{L^{\infty}(\mathbb{R}^{3})}\leq C^{\prime}\left\|u_{\epsilon}\right\|_{H_{r}^{1}(\mathbb{R}^{3})}\leq C^{\prime}M_{0}.$
Let $C^{\prime}M_{0}=m_{0}$. Hence, $\left\|u_{\epsilon}\right\|_{L^{\infty}(\mathbb{R}^{3})}\leq m_{0}$. The proof is completed. ∎
The proof of Theorem 1. According to Lemmas 1 and 2, for $0<\epsilon<\epsilon_{*}$, there exists a nontrivial weak solution $u$ satisfies
$\begin{cases}-\Delta u+V(x)u-\phi u=\frac{1}{\epsilon}\tilde{F}^{\prime}(\epsilon u)&\quad\text{in }\mathbb{R}^{3},\\ -\Delta\phi=u^{2}&\quad\text{in }\mathbb{R}^{3}.\end{cases}$
Furthermore, take $\epsilon^{*}=\min\{\frac{1}{m_{0}^{2}},\epsilon_{*}\}$. For $0<\epsilon<\epsilon^{*}$, it follows from (3.10) that
$\begin{split}\tilde{F}^{\prime}(\epsilon u)&=d^{\prime}(\epsilon u/\sqrt{\epsilon})F(\epsilon u)\frac{1}{\sqrt{\epsilon}}+d(\epsilon u/\sqrt{\epsilon})f(\epsilon u)-d^{\prime}(\epsilon u/\sqrt{\epsilon})\frac{1}{\sqrt{\epsilon}}|t|^{3}+3(1-d(\epsilon u/\sqrt{\epsilon}))|\epsilon u|\epsilon u\\ &=f(\epsilon u),~a.e.~\text{in}~x\in\mathbb{R}^{3},\end{split}$
where the fact $|\epsilon u/\sqrt{\epsilon}|\leq m_{0}\sqrt{\epsilon}\leq 1$ has been used.
[…]
4. Schrödinger-Poisson system with the sign-changing potential
The purpose of this section is to obtain a nontrivial solutions of system (1.1) with the sign-changing potential. In this case, the corresponding functional $I_{\epsilon}$ defined in (4.1) is different from (2.4). Before proving our results, we shall introduce the suitable function space where the critical points of the modified functional $I_{\epsilon}$ belong to.
Define the working space as follows:
$H^{1}_{V}(\mathbb{R}^{3})=\{u\in H^{1}(\mathbb{R}^{3}):~\int_{\mathbb{R}^{3}}V(x)u^{2}\text{d}x<\infty\},$
where the inner product
$\langle u,v\rangle=\int_{\mathbb{R}^{3}}\left(\nabla u\cdot\nabla v+\tilde{V}(x)uv\right)\text{d}x$
with $\tilde{V}$ defined in (1.5). Then the corresponding norm is $\|u\|_{H^{1}_{V}(\mathbb{R}^{3})}=\langle u,u\rangle^{\frac{1}{2}}$.
If $(V_{1})$ holds, the embedding $H^{1}_{V}(\mathbb{R}^{3})\hookrightarrow L^{s}(\mathbb{R}^{3})$ is compact for any $s\in[2,6)$ (see Bartsch and Wang [5]). Let $-\infty<\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{n}\leq\cdots$ be the complete sequence of eigenvalues of
$\Delta u+V(x)u=\lambda u,\ u\in H^{1}_{V}(\mathbb{R}^{3}).$
Each eigenvalue is repeated according to its multiplicity, and let $e_{1},e_{2},\cdots$ be the corresponding orthonormal eigenfunctions in $L^{2}(\mathbb{R}^{N})$.
4.1. Cerami condition
In order to prove Theorem 2, we firstly need to show that the functional $I_{\epsilon}$ satisfies the Cerami conditions. To this end, we rewrite $I_{\epsilon}$ in the following form
(4.1) $I_{\epsilon}(u)=\frac{1}{2}\int_{\mathbb{R}^{3}}\left(|\nabla u|^{2}+\tilde{V}(x)u^{2}\right)\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\text{d}x-\frac{1}{\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}(\epsilon u)\text{d}x-\frac{m}{2}\int_{\mathbb{R}^{3}}u^{2}\text{d}x.$
Lemma 4.
Let $\{u_{n}\}_{n=1}^{\infty}\subset H^{1}_{V}(\mathbb{R}^{3})$ be a Cerami sequence for the functional $I_{\epsilon}$. There exists a constant $\epsilon_{**}>0$ such that if $\epsilon\in(0,\epsilon_{**})$, then $u_{n}\rightarrow u\text{ in }H^{1}_{V}(\mathbb{R}^{3})$ with $u\in H^{1}_{V}(\mathbb{R}^{3})$.
Proof.
Let $\{u_{n}\}$ be a Cerami sequence of $I_{\epsilon}$, that is, for some $c\in\mathbb{R}$,
$I_{\epsilon}(u_{n})\to c~\text{and}~(1+\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})})I_{\epsilon}^{\prime}(u_{n})\to 0\ ~\text{as}~n\to+\infty.$
Firstly, we claim that
$\rho_{n}:=\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}<+\infty.$
If it is not true, we may assume $\rho_{n}\to+\infty$. Let $v_{n}=\frac{u_{n}}{\rho_{n}}$. Consequently, $\{v_{n}\}$ is bounded in $H^{1}_{V}(\mathbb{R}^{3})$. Up to a subsequence, $H^{1}_{V}(\mathbb{R}^{3})$ is compactly embedded in $L^{2}(\mathbb{R}^{3})$, we may assume that
$\begin{split}v_{n}&\to v~\text{weakly in}~H^{1}_{V}(\mathbb{R}^{3});\\ v_{n}&\to v~\text{strongly in}~L^{2}(\mathbb{R}^{3});\\ v_{n}&\to v~\text{a.e. in}~\mathbb{R}^{3}.\end{split}$
Consider
$\begin{split}c+o_{n}(1)&=I_{\epsilon}(u_{n})-\frac{1}{3}\langle I_{\epsilon}^{\prime}(u_{n}),u_{n}\rangle\\ &=\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+\tilde{V}(x)u_{n}^{2}\right)\text{d}x+\left(\frac{1}{3}-\frac{1}{4}\right)\int_{\mathbb{R}^{3}}\phi_{u_{n}}u_{n}^{2}\text{d}x-\frac{1}{\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}(\epsilon u_{n})\text{d}x\\ &\quad\quad-\frac{1}{3\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}^{\prime}(\epsilon u_{n})\epsilon u_{n}\text{d}x-\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}mu^{2}_{n}\text{d}x\\ &\geq\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+V(x)u_{n}^{2}\right)\text{d}x+\left(\frac{1}{3}-\frac{1}{4}\right)\int_{\mathbb{R}^{3}}\phi_{u_{n}}u_{n}^{2}\text{d}x\\ &\quad\quad-\left(1+\frac{1}{3}\right)\int_{\mathbb{R}^{3}}\left(\delta|u_{n}|^{2}+\epsilon C_{\delta}|u_{n}|^{3}\right)\text{d}x-\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}mu^{2}_{n}\text{d}x\\ &\geq\left(\left(\frac{1}{2}-\frac{1}{3}\right)-\frac{1}{4}\left(\frac{1}{3}-\frac{1}{4}\right)-\delta\left(1+\frac{1}{3}\right)\right)\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+V(x)u_{n}^{2}\right)\text{d}x\\ &\quad\quad+\left(\frac{1}{3}-\frac{1}{4}-C_{\delta}\epsilon\right)\int_{\mathbb{R}^{3}}|u_{n}|^{3}\text{d}x-\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}mu^{2}_{n}\text{d}x.\end{split}$
Take $\delta=\frac{1}{64}$. Then, there exists a constant $\epsilon_{*}>0$ small enough such that when $0<\epsilon\leq\epsilon_{*}$, one has
$\left(\frac{1}{3}-\frac{1}{4}-C_{\frac{1}{64}}\epsilon\right)\geq 0$
and then
$c+o_{n}(1)\geq\frac{1}{8}\|u_{n}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}-\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}mu^{2}_{n}\text{d}x.$
Multiplying both sides by $\rho_{n}^{-2}$, for large $n$, we have
$\left(\frac{1}{2}-\frac{1}{3}\right)m\|v_{n}\|_{L^{2}(\mathbb{R}^{3})}^{2}\geq\frac{1}{16}.$
Note that $v_{n}\to v$ strongly in $L^{2}(\mathbb{R}^{3})$. Then $v\neq 0$. Thus the Lebesgue measure of the set
$\mathcal{M}=\{x\in\mathbb{R}^{3}:v(x)\neq 0\}$
is positive. For $x\in\mathcal{M}$, as $n\to+\infty$, we deduce
$v_{n}(x)\to v(x)\neq 0~\text{and}~u_{n}(x)=v_{n}(x)\rho_{n}\to+\infty.$
On the other hand, it follows from assumptions $({\bf f})$ and $({\bf f^{\prime}})$ that
(4.2) $\tilde{F}(t)\geq 0\ ~\text{for any}~\epsilon\in(0,\frac{\delta_{0}^{2}}{4}).$
From (4.2) and the Fatou lemma, we have
$\begin{split}&\indent\int_{\mathbb{R}^{3}}\left(\frac{\phi_{u_{n}}u^{2}_{n}}{4}+\frac{1}{\epsilon^{2}}\tilde{F}(\epsilon u_{n})+\frac{1}{2}mu_{n}^{2}\right)\rho^{-2}_{n}\text{d}x\\ &=\int_{\mathbb{R}^{3}}\left(\frac{\phi_{u_{n}}u^{2}_{n}}{4u_{n}^{2}}+\frac{\tilde{F}(\epsilon u_{n})}{\epsilon^{2}u_{n}^{2}}+\frac{mu_{n}^{2}}{2u_{n}^{2}}\right)v^{2}_{n}\text{d}x\\ &\geq\frac{1}{4}\int_{\mathcal{M}}\phi_{u_{n}}v^{2}_{n}\text{d}x\to+\infty\end{split}$
as $n\to\infty$. Thus
$\begin{split}c-1&\leq I_{\epsilon}(u_{n})\\ &=\frac{1}{2}\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+\tilde{V}(x)u_{n}^{2}\right)\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{u_{n}}u_{n}^{2}\text{d}x\\ &\indent-\frac{1}{\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}(\epsilon u_{n})\text{d}x-\frac{m}{2}\int_{\mathbb{R}^{3}}u_{n}^{2}\text{d}x\\ &\leq\rho_{n}^{2}\left(\frac{1}{2}-\int_{\mathbb{R}^{3}}\left(\frac{\phi_{u_{n}}u^{2}_{n}}{4}+\frac{1}{\epsilon^{2}}\tilde{F}(\epsilon u_{n})+\frac{1}{2}mu_{n}^{2}\right)\rho^{-2}_{n}(x)\text{d}x\right)\to$-$\infty\end{split}$
as $n\to\infty$. This is impossible. Therefore, our claim is true.
From the boundedness of $\{u_{n}\}$ and $H^{1}_{V}(\mathbb{R}^{3})$ is compactly embedded in $L^{p}(\mathbb{R}^{3})(p\in[2,6))$, up to a sub-sequence we may assume
$\begin{split}u_{n}&\to u~\text{weakly in}~H^{1}_{V}(\mathbb{R}^{3});\\ u_{n}&\to u~\text{strongly in}~L^{p}(\mathbb{R}^{3})(p\in[2,6));\\ u_{n}&\to u~\text{a.e. in}~\mathbb{R}^{3}.\end{split}$
Next let $\epsilon_{**}=\min\{\epsilon_{*},\frac{\delta^{2}_{0}}{4}\}$, we can follow the same argument as Lemma 2 to show that $\{u_{n}\}$ has a convergent subsequence. ∎
By Morse theory, we know that if $I$ satisfies the Cerami conditions, $u={\bf 0}$ is an isolated critical point of $I$ and the critical points set $\mathcal{K}:=\{u\in H^{1}_{V}(\mathbb{R}^{3}):I^{\prime}(u)=0\}=\{\bf 0\}$, then $C_{q}(I,\infty)\cong C_{q}(I,0)$ for all $q\in\mathbb{N}$. If $C_{q}(I,\infty)\ncong C_{q}(I,0)$ for some $q\in\mathbb{N}$ then $I$ must have a nontrivial critical point. Therefore, one has to compute these groups to get the nontrivial critical point.
Proposition 1.
(see Bartsch and Li [4]) Suppose that $I\in C^{1}(H^{1}_{V}(\mathbb{R}^{3}),\mathbb{R})$ satisfies the Cerami condition and $C_{k}(I,0)\neq C_{k}(I,\infty)$ for some $k\in\mathbb{N}$, then $I$ has a nontrivial critical point.
To compute the critical groups, we need to decompose the workspace $H^{1}_{V}(\mathbb{R}^{3})$. Since $0$ is not an eigenvalue of $-\Delta+V(x)$, we can assume that there exists an integer $d\geq 0$ such that $0\in(\lambda_{d},\lambda_{d+1})$, where we set $\lambda_{0}=$-$\infty$. For $d\geq 1$, let $H^{1}_{V}(\mathbb{R}^{3})^{-}=\mbox{span}\{e_{1},\cdots,e_{d}\}\ \mbox{and}\ H^{1}_{V}(\mathbb{R}^{3})^{+}=(H^{1}_{V}(\mathbb{R}^{3})^{-})^{\perp}.$ Then $H^{1}_{V}(\mathbb{R}^{3})^{-}$ and $H^{1}_{V}(\mathbb{R}^{3})^{+}$ are the negative space and positive space of the quadratic form $B(u)=\frac{1}{2}\int_{\mathbb{R}^{3}}\left(|\nabla u|^{2}+V(x)u^{2}\right)\text{d}x,$ respectively. Thus, there exits a constant $\eta>0$, such that
(4.3) $\pm B(u)\geq\eta\|u\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2},\ u\in H^{1}_{V}(\mathbb{R}^{3})^{\pm}.$
4.2. Critical groups at zero
In this section, we will use the following proposition to compute the critical groups of $I_{\epsilon}$ at zero.
Proposition 2.
(see Liu [14]) Suppose $I\in C^{1}(H^{1}_{V}(\mathbb{R}^{3}),\mathbb{R})$ has a local linking at zero with respect to the decomposition $H^{1}_{V}(\mathbb{R}^{3})=H^{1}_{V}(\mathbb{R}^{3})^{-}\oplus H^{1}_{V}(\mathbb{R}^{3})^{+}$, i.e. for some $r>0$.
(4.4) $I(u)\leq 0\ \mbox{for}\ u\in H^{1}_{V}(\mathbb{R}^{3})^{-}\cap B_{r},\ \mbox{and}\ I(u)>0\ \mbox{for}\ u\in(H^{1}_{V}(\mathbb{R}^{3})^{+}\setminus\{{\bf 0}\})\cap B_{r},$
where $B_{r}=\{u\in H^{1}_{V}(\mathbb{R}^{3}):\|u\|_{H^{1}_{V}(\mathbb{R}^{3})}<r\}$. If $d=\dim H^{1}_{V}(\mathbb{R}^{3})^{-}<\infty$, then $C_{d}(I,0)\neq 0$.
Lemma 5.
Under assumptions ${\bf(V_{1})}$, ${\bf(V_{2})}$, ${\bf(f)}$ and ${\bf(f^{\prime})}$. Then the functional $I_{\epsilon}$ has a local linking at zero with respect to decomposition $H^{1}_{V}(\mathbb{R}^{3})=H^{1}_{V}(\mathbb{R}^{3})^{-}\oplus H^{1}_{V}(\mathbb{R}^{3})^{+}$.
Proof.
Recall (2.2), for any $\delta>0$ and $\epsilon\in(0,\frac{\delta^{2}_{0}}{4})$, there is a constant $C_{\delta}$ such that
$0\leq\tilde{F}(t)\leq\delta t^{2}+C_{\delta}|t|^{3},$
where $C_{\delta}$ is a constant not depending on $\epsilon$. Hence, as $\|u\|_{H^{1}_{V}(\mathbb{R}^{3})}\to 0$, we get
$\int_{\mathbb{R}^{3}}\tilde{F}(u)\text{d}x\leq o(\|u\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2}).$
Thus for $u\in H^{1}_{V}(\mathbb{R}^{3})^{+}\setminus\{\bf 0\}$ and $\|u\|_{H^{1}_{V}(\mathbb{R}^{3})}$ small enough, by (4.3) we obtain
$I_{\epsilon}(u)$ $=\frac{1}{2}\int_{\mathbb{R}^{3}}\left(|\nabla u|^{2}+V(x)u^{2}\right)\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\text{d}x-\frac{1}{\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}(\epsilon u)\text{d}x$
$\geq\frac{\eta}{4}\|u\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}-\frac{C}{4}\|u\|^{4}_{H^{1}_{V}(\mathbb{R}^{3})}+o(\|u\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2})$
$\geq\frac{\eta}{8}\|u\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}.$
Then for $r>0$ small enough, we obtain
(4.5) $I_{\epsilon}(u)>0\ \mbox{for}\ u\in(H^{1}_{V}(\mathbb{R}^{3})^{+}\setminus\{{\bf 0}\})\cap B_{r}.$
On the other hand, for $u\in H^{1}_{V}(\mathbb{R}^{3})^{-}\cap B_{r}$, it is easy to see that
$I_{\epsilon}(u)$ $=\frac{1}{2}\int_{\mathbb{R}^{3}}\left(|\nabla u|^{2}+V(x)u^{2}\right)\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\text{d}x-\frac{1}{\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}(\epsilon u)\text{d}x$
$\leq-\eta\|u\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\text{d}x-\frac{1}{\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}(\epsilon u)\text{d}x$
$\leq-\eta\|u\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}.$
From this and (4.5), it implies $I_{\epsilon}$ has a local linking property at zero. ∎
4.3. Critical groups at infinity
Lemma 6.
If assumptions ${\bf(V_{1})}$, ${\bf(V_{2})}$, ${\bf(f)}$ and ${\bf(f^{\prime})}$ hold. Then $C_{q}(I_{\epsilon},\infty)\cong 0$ with $q\in\mathbb{N}=\{0,1,2,\cdots\}$.
Proof.
Define $S^{\infty}$ as the unit sphere in $H^{1}_{V}(\mathbb{R}^{3})$. Firstly, we can show that
(4.6) $I_{\epsilon}(sh)\to-\infty\ \mbox{as}\ s\to+\infty\ \mbox{for any}\ h\in S^{\infty}.$
Similar with Lemma 4, from $\tilde{F}(t)\geq 0$ we deduce
$\begin{split}I_{\epsilon}(sh)&=\frac{1}{2}\int_{\mathbb{R}^{3}}\left(|\nabla(sh)|^{2}+\tilde{V}(x)(sh)^{2}\right)\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{(sh)}(sh)^{2}\text{d}x\\ &\indent-\frac{1}{\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}(\epsilon sh)\text{d}x-\frac{m}{2}\int_{\mathbb{R}^{3}}(sh)^{2}\text{d}x\\ &\leq s^{2}\left(\frac{1}{2}-\frac{s^{2}}{4}\int_{\mathbb{R}^{3}}\phi_{h}h^{2}\text{d}x\right)\to-\infty.\end{split}$
We next shall show the following claim is right.
Claim: There is a constant $A>0$ such that if $I_{\epsilon}(u)\leq-A$, then
$\frac{\mbox{d}}{\mbox{d}t}|_{t=1}I_{\epsilon}(tu)<0.$
If this claim is false, there exists a sequence $\{u_{n}\}\subset H^{1}_{V}(\mathbb{R}^{3})$ such that
(4.7) $I_{\epsilon}(u_{n})\leq-n\ \mbox{and}\ \langle I_{\epsilon}^{\prime}(u_{n}),u_{n}\rangle=\frac{\mbox{d}}{\mbox{d}t}|_{t=1}I_{\epsilon}(tu_{n})\geq 0.$
From (4.7), similar with (3.2), we have
(4.8) $\begin{split}0&\geq I_{\epsilon}(u_{n})-\frac{1}{3}\langle I_{\epsilon}^{\prime}(u_{n}),u_{n}\rangle\\ &=\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+\tilde{V}(x)u_{n}^{2}\right)\text{d}x+\left(\frac{1}{3}-\frac{1}{4}\right)\int_{\mathbb{R}^{3}}\phi_{u_{n}}u_{n}^{2}\text{d}x-\frac{1}{\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}(\epsilon u_{n})\text{d}x\\ &\quad\quad-\frac{1}{3\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}^{\prime}(\epsilon u_{n})\epsilon u_{n}\text{d}x-\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}mu^{2}_{n}\text{d}x\\ &\geq\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+V(x)u_{n}^{2}\right)\text{d}x+\left(\frac{1}{3}-\frac{1}{4}\right)\int_{\mathbb{R}^{3}}\phi_{u_{n}}u_{n}^{2}\text{d}x\\ &\quad\quad-\left(1+\frac{1}{3}\right)\int_{\mathbb{R}^{3}}\left(\delta|u_{n}|^{2}+\epsilon C_{\delta}|u_{n}|^{3}\right)\text{d}x-\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}mu^{2}_{n}\text{d}x\\ &\geq\frac{1}{8}\|u_{n}\|^{2}_{H^{1}_{V}(\mathbb{R}^{3})}-\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}mu^{2}_{n}\text{d}x,\end{split}$
where $\delta$ denoted by $\frac{1}{64}$ and $\epsilon>0$ small.
Notice that if $\{u_{n}\}_{n=1}^{\infty}$ is a bounded sequence in $H^{1}_{V}(\mathbb{R}^{3})$, then $\{I_{\epsilon}(u_{n})\}$ is also bounded. Then by $I_{\epsilon}(u_{n})\leq-n$, it implies
$\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+\widetilde{V}(x)|u_{n}|^{2}\right)\text{d}x\to+\infty\ \mbox{as}\ n\to+\infty.$
Setting $v_{n}=\frac{u_{n}}{\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}}$, $\{v_{n}\}_{n=1}^{\infty}$ is a bounded sequence in $H^{1}_{V}(\mathbb{R}^{3})$. Up to subsequence, one has
$v_{n}\to v\ \mbox{ weakly in}\ H^{1}_{V}(\mathbb{R}^{3}),\ v_{n}\to v\ \mbox{ strongly in}\ L^{2}(\mathbb{R}^{3}),\ v_{n}\to v\ \mbox{a.e. in}\ \mathbb{R}^{3}.$
Multiplying both side of (4.8) by $\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}$, we obtain
$\left(\frac{1}{2}-\frac{1}{3}\right)\int_{\mathbb{R}^{3}}mv^{2}_{n}\text{d}x\geq\frac{1}{8}.$
From $v_{n}\to v$ strongly in $L^{2}(\mathbb{R}^{3})$, it follows that $v\neq 0$. Thus the Lebesgue measure of the set
$\mathcal{M}=\{x\in\mathbb{R}^{3}:v(x)\neq 0\}$
is positive.
By the definition of $\tilde{F}(t)$, conditions $({\bf f})$ and $({\bf f^{\prime}})$, for any $\epsilon\in(0,\frac{\delta_{0}}{4})$, it follows that
(4.9) $\tilde{F}^{\prime}(t)t\geq 0.$
For $x\in\mathcal{M}$, as $n\to+\infty$, we deduce
$v_{n}(x)\to v(x)\neq 0~\text{and}~u_{n}(x)=v_{n}(x)\rho_{n}\to+\infty.$
It follows from (4.9) and the Fatou lemma that
$\begin{split}&\indent\int_{\mathbb{R}^{3}}\left(\phi_{u_{n}}u^{2}_{n}+\frac{1}{\epsilon}\tilde{F}^{\prime}(\epsilon u_{n})u_{n}+mu_{n}^{2}\right)\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{-2}\text{d}x\\ &=\int_{\mathbb{R}^{3}}\left(\frac{\phi_{u_{n}}u^{2}_{n}}{u_{n}^{2}}+\frac{\tilde{F}^{\prime}(\epsilon u_{n})\epsilon u_{n}}{\epsilon^{2}u_{n}^{2}}+\frac{mu_{n}^{2}}{u_{n}^{2}}\right)v^{2}_{n}\text{d}x\\ &\geq\int_{\mathcal{M}}\phi_{u_{n}}v^{2}_{n}\text{d}x\to+\infty\end{split}$
as $n\to\infty$. Thus
$\begin{split}0&\leq\langle I_{\epsilon}^{\prime}(u_{n}),u_{n}\rangle\\ &=\int_{\mathbb{R}^{3}}\left(|\nabla u_{n}|^{2}+\tilde{V}(x)u_{n}^{2}\right)\text{d}x-\int_{\mathbb{R}^{3}}\phi_{u_{n}}u_{n}^{2}\text{d}x\\ &\indent-\frac{1}{\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}^{\prime}(\epsilon u_{n})\epsilon u_{n}\text{d}x-m\int_{\mathbb{R}^{3}}u_{n}^{2}\text{d}x\\ &\leq\|u_{n}\|_{H^{1}_{V}(\mathbb{R}^{3})}^{2}\left(\frac{1}{2}-\int_{\mathcal{M}}\phi_{u_{n}}v^{2}_{n}\text{d}x\right)\to$-$\infty\end{split}$
as $n\to\infty$. This is impossible. Therefore, our claim is true.
Due to this claim and (4.6), for $B>A$ large enough, there exists an unique $T:=T(u)>0$ such that
$I_{\epsilon}(T(u)u)=-B\ \ \text{for every }u\in S^{\infty}.$
It follows from the implicit function theorem that
$T$ is a continuous function from $S^{\infty}$ to $\mathbb{R}$.
Thus the deformation retract $\eta:[0,1]\times(H^{1}_{V}(\mathbb{R}^{3})\setminus B_{1})\to H^{1}_{V}(\mathbb{R}^{3})$ denoted as
$\eta(s,u)=(1-s)u+sT(u)u$
with $\eta(0,u)=u$, $\eta(1,u)\in I_{\epsilon}^{-B}$. Thus
$C_{q}(I_{\epsilon},\infty)=H^{q}(H^{1}_{V}(\mathbb{R}^{3}),I_{\epsilon}^{-B})\cong H^{q}(H^{1}_{V}(\mathbb{R}^{3}),H^{1}_{V}(\mathbb{R}^{3})\setminus B_{1})\cong 0,\ \mbox{for all}\ q\in\mathbb{N}.$
This completes the proof of Lemma 6. ∎
[…]
5. The proof of Theorem 3
Now we study the asymptotical behavior of non-trivial solutions in Theorem 1.
[…]
Let
$I_{0}(u)=\frac{1}{2}\int_{\mathbb{R}^{3}}\left(|\nabla u|^{2}+V(x)u^{2}\right)\text{d}x-\frac{1}{4}\int_{\mathbb{R}^{3}}\phi_{u}u^{2}\text{d}x$
be the energy functional of the following Choquard equation
(5.1) $\Delta u+V(x)u=\frac{1}{4\pi}\int_{\mathbb{R}^{3}}\frac{u^{2}(y)}{|x-y|}\text{d}y\ \ ~\text{in}~\mathbb{R}^{3}.$
Then it is easy to say $I_{0}\in C^{1}$ in $H^{1}_{r}(\mathbb{R}^{3})$. Since $H^{1}_{r}(\mathbb{R}^{3})$ is compactly embedded in $L^{p}(\mathbb{R}^{3})(2<p<6)$, $I_{0}$ satisfies the (PS) condition.
By (2.2), one has
$\begin{split}\lim\limits_{\epsilon\to 0}\left(I_{\epsilon}(u_{\epsilon})-I_{0}(u_{\epsilon})\right)=\lim\limits_{\epsilon\to 0}\frac{1}{\epsilon^{2}}\int_{\mathbb{R}^{3}}\tilde{F}(\epsilon u_{\epsilon})\text{d}x=0.\end{split}$
Similarly, for any $v\in H^{1}_{r}(\mathbb{R}^{3})$, it implies
$\begin{split}\lim\limits_{\epsilon\to 0}\langle I_{\epsilon}^{\prime}(u_{\epsilon})-I_{0}^{\prime}(u_{\epsilon}),v\rangle=\lim\limits_{\epsilon\to 0}\frac{1}{\epsilon}\int_{\mathbb{R}^{3}}\tilde{f}(\epsilon u_{\epsilon})v\text{d}x=0.\end{split}$
Thus $\{u_{\epsilon}\}$ is a (PS) sequence of $I_{0}$. Then by the (PS) condition of $I_{0}$, we have
$u_{\epsilon}\to u_{0}~\text{strongly~in}~H^{1}_{r}(\mathbb{R}^{3})\ ~\text{as}~\epsilon\to 0,$
where $u_{0}$ is a critical point of $I_{0}$ and it is also a weak solution of equation (5.1). Recall that $\lambda=\frac{1}{\epsilon^{2}}$ and $u^{\lambda}=\epsilon u_{\epsilon}$, hence Theorem 3 has been proved.
Acknowledgement. S. Liang was supported by the Young Outstanding Talents Project of Scientific Innovation and entrepreneurship in Jilin (No. 20240601048RC), the National Natural Science Foundation of China (No. 12371455). L. Ma was supported by National Natural Science Foundation of China (No. 12301288).